A 6 V battery is connected across three 100 ohm resistors in series. What is the current through each resistor?


1
Expert's answer
2022-01-02T18:23:05-0500

Let's first find the equivalent resistance of three 100 ohm resistors connected in series:


"R_{eq}=R_1+R_2+R_3,""R_{eq}=100\ \Omega+100\ \Omega+100\ \Omega=300\ \Omega."

Since the current through each resistor connected in series is the same, we can write:


"I=I_1=I_2=I_3=\dfrac{V}{R_{eq}}=\dfrac{6\ V}{300\ \Omega}=0.02\ A."